OpenStudy (anonymous):

Use basic identities to simplify the expression. one divided by cotangent of theta to the second power. + sec θ cos θ

OpenStudy (saurabh135):

did u mean \[ 1/\cot ^{2} \theta + \sec \theta*\cos \theta\]